OpenStudy (pj754):

How do you isolate t when y=2sin(t)? How do I get rid of the sin?

OpenStudy (pj754):

y/2=sin(t), how do I go from there?

sam (.sam.):

Example\[\sin(30)=\frac{1}{2} \\ \\ \sin^{-1}(\frac{1}{2})=30\]

sam (.sam.):

\[\sin(t)=\frac{y}{2} \\ \\ t=\sin^{-1}(?)\]

OpenStudy (pj754):

oh so arcsin(y/2)=t right?

sam (.sam.):

Mhm

OpenStudy (pj754):

thanks!

sam (.sam.):

No problem